OpenStudy (anonymous):

What is the value of h(–2) when h(x) = 3x + 1? Please help

OpenStudy (anonymous):

Or can someone explain to me how functions work?

OpenStudy (anonymous):

substitute -2 for x so 3(-2)+1

OpenStudy (anonymous):

So basically, you replace the value of x in your function by -2. h(-2)=3(-2)+1 h(-2)=-6+1 h(-2)=-5

OpenStudy (anonymous):

ok, so the answer would be -5?